OpenStudy (anonymous):

So, on my right triangle triangle I have here, they give all the sides and are asking for an angle The sides are Opp=15 Adj=8 Hyp=17 And we are looking for the x angle.

OpenStudy (anonymous):

Do you know SOHCAHTOA? Your sine, cosine, and tangent laws?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Okay, since you know all three side lengths, you can use any of the three functions to determine x.

OpenStudy (anonymous):

For example, \(\sin x = \frac{\text{opp}}{\text{hyp}}\)

OpenStudy (anonymous):

And that is it? Just use 17 over 15 ,divide and my answer would be 1.13?

OpenStudy (anonymous):

It's not x = 1.13. It's sin x = 1.13.

OpenStudy (anonymous):

Ah, sorry for not using correct mathematical vocabulary.

OpenStudy (anonymous):

No, that's not my point. The point is, you need to solve for x. Right now, you only know sin x.

OpenStudy (anonymous):

You need to get rid of the sin.

OpenStudy (anonymous):

Oh? How is this done?

OpenStudy (anonymous):

One thing first. It's not 17/15, it's 15/17.

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

So, \(\sin x = \frac{15}{17} = 0.882\)

OpenStudy (anonymous):

To get rid of the sin, you calculate \(\sin^{-1}\). On your calculator, you should have a way to do this.

OpenStudy (anonymous):

So, \(x = \sin^{-1} (\frac{15}{17})\)

OpenStudy (anonymous):

I did not know there was such an option? Thank you.

OpenStudy (anonymous):

You might need to hit something like 2nd function, or inverse, then hit sin.

OpenStudy (anonymous):

You should end up with something around 60 degrees, if it's sine of 15/17.

OpenStudy (anonymous):

This isn't my outcome, I get a decimal of 1.001

OpenStudy (anonymous):

Okay, so you took your amount (0.882) and hit inverse sine (sin^-1) and got 1.001?

OpenStudy (anonymous):

Oh but wait, that is sin(h), Sorry, personal mistake.

OpenStudy (anonymous):

No worries.

OpenStudy (anonymous):

Since you knew all the sides, you could have also done \(x = \cos^{-1} (\frac{8}{17})\) or \(x = \tan^{-1} (\frac{15}{8})\), and gotten identical answers.

OpenStudy (anonymous):

Odd, I get the decimal of 0.795, how does this lead to 60 degrees?

OpenStudy (anonymous):

What calculation gives you 0.795?

OpenStudy (anonymous):

Well, seeing as I'm using the computer based calculator, this may have swayed my calculation. When typed in, it shows up as *asinh(15 /17)*

OpenStudy (anonymous):

Where did you type that in?

OpenStudy (anonymous):

But lo, I have found my answer as you have predicted, thank you so much kind sir, It was simply a typing error.

OpenStudy (anonymous):

Good stuff! You're very welcome. :)

OpenStudy (anonymous):

I shall be on my way! Good day to you, and thank you for your time.

OpenStudy (anonymous):

You too, David! Cheers!